OpenStudy (anonymous):

Find the area of the region that lies inside the first curve and outside the second curve. r = 11 cos θ, r = 5 + cos θ

OpenStudy (irishboy123):

 \[A = 2 \int\limits_{\theta = 0}^{\pi/3} \space{} \int\limits_{r = 5 + \cos \theta}^{11 \cos \theta} r dr d \theta\]